# CMSO-transducing tree-like graph decompositions

### v2

@@ 4.1 Transducing weakly-partitive trees
If $\mathcal{S}$ is partitive then the weakly-partitive tree it induces enjoys a simple form, and is truly unique. Indeed, the label $\mathsf{linear}$ and, thus, the partial order $<$, are not needed.
Theorem 4.6 ([7]).
Let $\mathcal{S}$ be a weakly-partitive family and $(T,\lambda,<)$ be the weakly-partitive tree it induces. If $\mathcal{S}$ is partitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and $<$ is empty.
Hence, in case of a partitive set systems $(U,\mathcal{S})$, we can consider the simpler object $(T,\lambda)$, called the partitive tree induced by $\mathcal{S}$ (or the partitive tree of $(U,\mathcal{S})$) in which $\lambda$ maps $V(T)\setminus L(T)$ to $\{\mathsf{degenerate},\mathsf{prime}\}$. As a direct consequence of Theorems 4.6 and 4.4, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the partitive tree induced by a partitive set system and naturally modeled by an $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure.
Corollary 4.7.
@@
4.2 Application to modular decomposition
@@
Let $G$ be a directed graph and let $M\subseteq V(G)$. We say that $M$ is a module (of $G$) if for every $u\notin M$ and every $v,w\in M$, ${uv\in E(G)\!\iff\!uw\in E(G)}$ and ${vu\in E(G)\!\iff\!wu\in E(G)}$. Clearly, the empty set, $V(G)$, and all the singletons $\{x\}$ for $x\in V(G)$ are modules; they are called the trivial modules of $G$. We say a non-empty module $M$ is maximal if it is not properly contained in any non-trivial module. Let $M$ and $M^{\prime}$ be two disjoint non-empty modules of $G$. Considering the edges that go from $M$ to $M^{\prime}$, namely edges from the set $(M\times M^{\prime})\cap E(G)$, we have two possibilities: either it is empty, or it is equal to $M\times M^{\prime}$. We write $M\text{$\not\to$}M^{\prime}$ in the former case and $M\text{$\to$}M^{\prime}$ in the latter. (It is of course possible to have both $M\text{$\to$}M^{\prime}$ and $M^{\prime}\text{$\to$}M$.) A modular partition of $G$ is a partition $\mathcal{P}=\{M_{1},\ldots,M_{\ell}\}$ of $V(G)$ such that every $M_{i}$ is a non-empty module. A modular partition $\mathcal{P}=\{M_{1},\dots,M_{\ell}\}$ is called maximal if it is non-trivial and every $M_{i}$ is strong and maximal. Note that every graph has exactly one maximal modular partition.
@@
A modular decomposition of $G$ is a rooted tree $T$ in which the leaves are the vertices of $G$, and for each inner node $t\in T$, $t$ has at least two children and the set $L(T_{t})$ is a module of $G$. In a modular decomposition $T$ of $G$, for each inner node $t\in V(T)$ with children $c_{1},\ldots,c_{r}$, the family $\mathcal{P}_{t}=\{L(T_{c_{1}}),\ldots,L(T_{c_{r}})\}$ is a modular partition of $G[V(T_{t})]$. When each such partition is maximal, the decomposition is unique and it is called the maximal modular decomposition of $G$. The maximal modular decomposition $T$ of $G$ alone is not sufficient to characterize $G$. However, enriching $T$ with, for each inner node $t$ with children $c_{1},\ldots,c_{j}$, the information of which pair of modules $\big{(}L(T_{c_{i}}),L(T_{c_{j}})\big{)}$ is such that $L(T_{c_{i}})\text{$\to$}L(T_{c_{j}})$, yields a unique canonical representation of $G$. Formally, the enriched modular decomposition of $G$ is the pair $(T,F)$ where $T$ is the maximal modular decomposition of $G$ (with $L(T)=V(G)$) and $F\subset V(T)\times V(T)$ is a binary relation, that relates a pair $(s,t)$ of nodes of $T$, denoted $st\in F$, exactly when $s$ and $t$ are siblings and $L(T_{s})\text{$\to$}L(T_{t})$. The elements of $F$ are called $\mathsf{m}$-edges.
@@
It should be mentioned that the family of all non-empty modules of $G$ is known to be weakly-partitive (or even partitive when $G$ is undirected). In particular, the maximal modular decomposition $T$ of $G$ is the laminar tree induced by the family of strong modules. Hence Theorem 4.4 could be used to produce a partially-ordered and labeled tree which displays all the modules of $G$. However, this weakly-partitive tree is not sufficient for being able to recover the graph $G$ from it. We now prove how to obtain the enriched modular decomposition of $G$ through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction.
@@
To model enriched modular decompositions as relational structures we use the relational vocabulary $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$ where $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ are two binary relation names. An enriched modular decomposition $(T,F)$ of a graph $G$ is modeled by the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ with universe $U_{\mathbb{M}}=V(T)$, $\mathsf{ancestor}_{\mathbb{M}}$ being the set of pairs $(s,t)$ for which $s$ is an ancestor of $t$ in $T$, and $\textsf{m}\text{-}\mathsf{edge}_{\mathbb{M}}$ being the set of all pairs $(s,t)$ such that $st\in F$ (in particular, $s$ and $t$ are siblings in $T$). We use Lemma 4.1 in order to transduce the maximal modular decomposition of a graph.
Theorem 4.8.
@@
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction given by Lemma 4.1 which produces the $\{\mathsf{edge},\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{S}$;
@@
In order to define $\tau_{1}$ it is sufficient to observe that there exists an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET}}(Z)$ with one monadic free-variable, which is satisfied exactly when $Z$ is a non-empty module of $G$. Then, since $\tau_{2}$ is given by Lemma 4.1, it only remains to define $\tau_{3}$. Given an inner node $t$, we can select, within $\operatorname{MSO}$, the set $L(T_{t})$ of leaves of the subtree rooted in $t$. We thus assume a function $\mathsf{leafset}$, with one first-order free-variable which returns the set of leaves of the subtree rooted at the given node. Equipped with this function, we can define the $\operatorname{MSO}$-formula $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ which selects pairs $(s,r)$ of siblings such that $sr\in F$. Remember that this happen exactly when there exist $u\in L(T_{s})$ and $v\in L(T_{r})$ such that $uv\in E(G)$. Hence, $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ could be defined as:
@@
Let $G$ be a graph, $(T,F)$ be its modular decomposition, and $(T,\lambda,<)$ be the weakly-partitive tree induced by the (weakly-partitive) family of its modules. Let $t$ be an inner node of $T$, let $\mathcal{C}$ be its set of children, and let $C$ be the graph $\big{(}V(T)\setminus L(T),F\big{)}\left[\mathcal{C}\right]$ induced by $F$ on the set of children of $t$. It can be checked that, if $\lambda(t)=\mathsf{degenerate}$ then $C$ is either a clique or an independant, and if $\lambda(t)=\mathsf{linear}$ then $C$ is a tournament consistent with $<_{t}$ (i.e., for every $x,y\in\mathcal{C}$, $xy$ is an edge of $C$ if and only if $x<_{t}y$) or with the inverse of $<_{t}$. In the former case, we can refine the $\mathsf{degenerate}$ label into $\mathsf{series}$ and $\mathsf{parallel}$ labels, thus expressing that $C$ is a clique or an indenpendant, respectively. In the latter case, up-to reversing $<_{t}$, we can ensure that $C$ is a tournament consistent with $<_{t}$. This yields a refined weakly-partitive tree $(T,\gamma,<)$, where $\gamma$ maps inner nodes to $\{\mathsf{series},\mathsf{parallel},\mathsf{prime},\mathsf{linear}\}$ and $<$ is the order $\bigcup_{t\in\gamma^{-1}(\mathsf{linear})}<_{t}$ which ensures that tournaments are consistent with the corresponding $<_{t}$. Notice that this labeled and partially-ordered tree is now uniquely determined from $G$. Moreover, edges from $F$ that connect children of a node not labeled by $\mathsf{prime}$ can be recovered from the so-refined weakly-partitive tree. In particular, if no nodes of $T$ is labeled by $\mathsf{prime}$, $(T,F)$ and thus $G$ is fully characterized by $(T,\gamma,<)$. Graphs for which this property holds are known as directed cographs, and can be described by the refined weakly-partitive tree, called cotree, explained above and formalized in the following statement.
Theorem 4.10.
@@
Let $G$ be a directed cograph and let $T$ be the laminar tree induced by the family of its strong modules. There exists a unique total labeling $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{series},\mathsf{parallel},\mathsf{linear}\}$ of labels, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a unique linear ordering $<_{t}$ of its children, such that every inner node having exactly two children is labeled $\mathsf{series}$ or $\mathsf{parallel}$, and the following condition is satisfied:
@@
  • for every two leaves $x$ and $y$ of $T$, denoting by $t$ their least common ancestor, $xy$ is an edge of $G$ if and only if either $t$ is labeled by $\mathsf{linear}$ and $x<_{t}y$, or $t$ is labeled by $\mathsf{series}$.
@@
We naturally model cotrees as $\{\mathsf{ancestor},\mathsf{series},\varepsilon\text{-}\mathsf{ord}\}$-structures as follow. A cotree $(T,\gamma,<)$ is modeled by $\mathbb{C}$ where $U_{\mathbb{C}}=V(T)$, $\mathsf{series}_{\mathbb{C}}=\gamma^{-1}(\mathsf{series})$, and $\varepsilon\text{-}\mathsf{ord}_{\mathbb{C}}=\{(x,y)\mid x<y\}$. The nodes that are labeled by $\mathsf{linear}$ could be recovered as those inner nodes whose children are related by $\varepsilon\text{-}\mathsf{ord}$, while the nodes that are labeled by $\mathsf{parallel}$ could be recovered as those inner nodes which are labeled neither by $\mathsf{series}$ nor by $\mathsf{linear}$. Based on Theorem 4.10 and as a consequence of Theorem 4.8, we can design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which produces the cotree of a cograph $G$ from $G$.
Corollary 4.11.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each cograph $G$ modeled by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\mathsf{series},\varepsilon\text{-}\mathsf{ord}\}$-structure $\mathbb{C}$ representing the cotree $(T,\gamma,<)$ of $G$.
5 Transducing split and bi-join decompositions

### v4

@@ 4.1. Transducing weakly-partitive trees
If $\mathcal{S}$ is partitive then the weakly-partitive tree it induces enjoys a simple form, and is unique. Indeed, the label $\mathsf{linear}$ and, thus, the partial order <<, are not needed.{thmC}[[CHM81]] Let $\mathcal{S}$ be a weakly-partitive family and $(T,\lambda,<)$ be the weakly-partitive tree it induces. If $\mathcal{S}$ is partitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and << is empty. Hence, in case of a partitive set systems $(U,\mathcal{S})$, we can consider the simpler object $(T,\lambda)$, called the partitive tree induced by $\mathcal{S}$ (or the partitive tree of $(U,\mathcal{S})$) in which $\lambda$ maps $V(T)\setminus L(T)$ to $\{\mathsf{degenerate},\mathsf{prime}\}$. As a direct consequence of 4.2 and of 15, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the partitive tree induced by a partitive set system and naturally modeled by an $\{\mathsf{ancestor},\mathsf{degenerate}\}$-structure.
Corollary 16.
@@
4.2. Application to modular decomposition
@@
Let $G$ be a directed graph and let $M\subseteq V(G)$. We say that $M$ is a module (of $G$) if for every $u\notin M$ and every $v,w\in M$, $uv\in E(G)\!\iff\!uw\in E(G)$} and $vu\in E(G)\!\iff\!wu\in E(G)$}. Clearly, the empty set, $V(G)$, and all the singletons $\{x\}$ for $x\in V(G)$ are modules; they are called the trivial modules of $G$. We say a non-empty module $M$ is maximal if it is not properly contained in any non-trivial module. Furthermore, we use the notion of strong modules to coincide with the strong sets in the set system consisting of all modules of $G$. Let $M$ and $M^{\prime}$ be two disjoint non-empty modules of $G$. Considering the edges that go from $M$ to $M^{\prime}$, namely edges from the set $(M\times M^{\prime})\cap E(G)$, we have two possibilities: either it is empty, or it is equal to $M\times M^{\prime}$. We write $M\text{$\not\to$}M^{\prime}$ in the former case and $M\text{$\to$}M^{\prime}$ in the latter. (It is of course possible to have both $M\text{$\to$}M^{\prime}$ and $M^{\prime}\text{$\to$}M$.) A modular partition of $G$ is a partition $\mathcal{P}=\{M_{1},\ldots,M_{\ell}\}$ of $V(G)$ such that every $M_{i}$ is a non-empty module. A modular partition $\mathcal{P}=\{M_{1},\dots,M_{\ell}\}$ is called maximal if it is non-trivial and every $M_{i}$ is strong and maximal. Note that every graph has exactly one maximal modular partition.
@@
A modular decomposition of $G$ is a rooted tree $T$ in which the leaves are the vertices of $G$, and for each inner node $\in T, tt$ has at least two children and the set $L(T_{t})$ is a module of $G$. In a modular decomposition $T$ of $G$, for each inner node $t\in V(T)$ with children $c_{1},\ldots,c_{r}$, the family $\mathcal{P}_{t}=\{L(T_{c_{1}}),\ldots,L(T_{c_{r}})\}$ is a modular partition of $G[L(T_{t})]$. When each such partition is maximal, the decomposition is unique and it is called the maximal modular decomposition of $G$. The maximal modular decomposition $T$ of $G$ alone is not sufficient to characterize $G$. However, enriching $T$ with, for each inner node $t$ with children $c_{1},\ldots,c_{j}$, the information of which pair of modules $\big{(}L(T_{c_{i}}),L(T_{c_{j}})\big{)}$ is such that $L(T_{c_{i}})\text{$\to$}L(T_{c_{j}})$, yields a unique canonical representation of $G$. Formally, the enriched modular decomposition of $G$ is the pair $(T,F)$ where $T$ is the maximal modular decomposition of $G$ (with $L(T)=V(G)$) and $F\subset V(T)\times V(T)$ is a binary relation, that relates a pair $(s,t)$ of nodes of $T$, denoted $st\in F$, exactly when $s$ and $t$ are siblings and $L(T_{s})\text{$\to$}L(T_{t})$. The elements of $F$ are called $\mathsf{m}$-edges.
@@
It should be mentioned that the family of all non-empty modules of $G$ is known to be weakly-partitive (or even partitive when $G$ is undirected). In particular, the maximal modular decomposition $T$ of $G$ is the laminar tree induced by the family of strong modules. Hence, 15 could be used to produce a partially-ordered and labeled tree which displays all the modules of $G$. However, this weakly-partitive tree is not sufficient for being able to recover the graph $G$ from it. We now prove how to obtain the enriched modular decomposition of $G$ through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction.
@@
To model enriched modular decompositions as relational structures we use the relational vocabulary $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$ where $\mathsf{ancestor}$ and $\textsf{m}\text{-}\mathsf{edge}$ are two binary relation names. An enriched modular decomposition $(T,F)$ of a graph $G$ is modeled by the $\{\mathsf{ancestor},\textsf{m}\text{-}\mathsf{edge}\}$-structure $\mathbb{M}$ with universe $U_{\mathbb{M}}=V(T)$, $\mathsf{ancestor}_{\mathbb{M}}$ being the set of pairs $(s,t)$ for which $s$ is an ancestor of $t$ in $T$, and $\textsf{m}\text{-}\mathsf{edge}_{\mathbb{M}}$ being the set of all pairs $(s,t)$ such that $st\in F$ (in particular, $s$ and $t$ are siblings in $T$). We use 14 in order to transduce the maximal modular decomposition of a graph.
Theorem 17.
@@
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction given by 14 which produces the $\{\mathsf{edge},\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{G}\sqcup\mathbb{S}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{S}$;
@@
In order to define $\tau_{1}$ it is sufficient to observe that there exists an $\operatorname{MSO}$-formula $\phi_{\mathsf{SET}}(Z)$ with one monadic free-variable, which is satisfied exactly when $Z$ is a non-empty module of $G$. Then, since $\tau_{2}$ is given by 14, it only remains to define $\tau_{3}$. Given an inner node $t$, we can select, within $\operatorname{MSO}$, the set $L(T_{t})$ of leaves of the subtree rooted in $t$. We thus assume a function $\mathsf{leafset}$, with one first-order free-variable which returns the set of leaves of the subtree rooted at the given node. Equipped with this function, we can define the $\operatorname{MSO}$-formula $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ which selects pairs $(s,r)$ of siblings such that $sr\in F$. Remember that this happen exactly when there exist $u\in L(T_{s})$ and $v\in L(T_{r})$ such that $uv\in E(G)$. Hence, $\phi_{\textsf{m}\text{-}\mathsf{edge}}$ could be defined as:
@@
In the following we provide an example of a natural property that can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-defined easily using the modular decomposition and hence (though backwards translation) is $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-definable on graphs. {exa} We want to define the property of an undirected graph $G$ having an even number of modules in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$. Let $T$ be the modular decomposition of $G$ and $\mathcal{M}$ the set of modules of $G$. As $\mathcal{M}$ is partitive, we can understand $T$ as a partitive tree and let $\lambda:V(T)\rightarrow\{\mathsf{prime},\mathsf{degenerate}\}$ be the labeling of $T$ that exists due to 4.1 and 4.2. By 5.2 it holds that the number of modules of $G[L(T_{t})]$ for a node $t$ with children $s_{1},\dots,s_{\ell}$ is exactly the sum of modules of the $G[L(T_{s_{i}})]$ if $\lambda(t)=\mathsf{prime}$ and the product of the $G[L(T_{s_{i}})]$ if $\lambda(t)=\mathsf{degenerate}$. Hence, given a set $X$ of nodes of $T$ we can check whether $X$ is precisely the set of nodes $t$ for which $G[L(T_{t})]$ has an even number of modules in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$. For this we check for each node $t$ their labeling and either the parity of children of $t$ that are in $X$ (if $\lambda(t)=\mathsf{prime}$) or the existence of a child in $X$ (if $\lambda(t)=\mathsf{degenerate}$) and can conclude whether $t$ should be included in $X$ or not. Note that additionally leaves of $T$ are not included in $X$ as a graph with one vertex only contains one module. The number of modules of $G$ is even if and only if there is a set $X$ which consist of all nodes $t$ for which $G[L(T_{t})]$ is even and the root of $T$ is in $X$.
To define the sentence, we let $\mathbb{T}$ be the $\{\mathsf{ancestor},\mathsf{prime},\mathsf{degenerate}\}$-structure modeling $T$ equipped with $\lambda$. While we did not provide a transduction producing this particular structure, it is easy to modify the above transductions to obtain the desired structure. We further use auxiliary predicates $\mathsf{children}(Y,y)$ expressing that $Y$ is the set of children of $y$, $\mathsf{root}(y)$ expressing that $y$ is the root of $T$ and $\mathsf{leaf}(y)$ expressing that $y$ is a leaf of $T$, which are routine to implement. Hence, we can define a sentence $\phi$ that is satisfied by any graph that has an even number of modules as follows.
$\phi:=$ $\exists X\forall y\exists Y\Bigg{(}\mathsf{children}(Y,y)\land\Big{[}\mathsf{root}(y)\rightarrow y\in X\Big{]}\land\Big{[}\mathsf{leaf}(y)\rightarrow y\notin X\Big{]}\land$
$\Big{[}\Big{(}\big{(}\mathsf{prime}(y)\land\operatorname{C}_{2}(X\cap Y)\big{)}\lor\big{(}\mathsf{degenerate}(y)\land X\cap Y\neq\emptyset\big{)}\Big{)}\leftrightarrow y\in X\Big{]}.$
@@
Let $G$ be a graph, $(T,F)$ be its modular decomposition, and $(T,\lambda,<)$ be the weakly-partitive tree induced by the (weakly-partitive) family of its modules. Let $t$ be an inner node of $T$, let $\mathcal{C}$ be its set of children, and let $C$ be the graph $\big{(}V(T)\setminus L(T),F\big{)}\left[\mathcal{C}\right]$ induced by $F$ on the set of children of $t$. It can be checked that, if $\lambda(t)=\mathsf{degenerate}$ then $C$ is either a clique or an independant, and if $\lambda(t)=\mathsf{linear}$ then $C$ is a tournament consistent with $_{t}$ (i.e., for every $\in\mathcal{C}, xyxy$ is an edge of $C$ if and only if $x<_{t}y$) or with the inverse of $_{t}$. In the former case, we can refine the $\mathsf{degenerate}$ label into $\mathsf{series}$ and $\mathsf{parallel}$ labels, thus expressing that $C$ is a clique or an independent set, respectively. In the latter case, up-to reversing $_{t}$, we can ensure that $C$ is a tournament consistent with $_{t}$. This yields a refined weakly-partitive tree $(T,\gamma,<)$, where $\gamma$ maps inner nodes to $\{\mathsf{series},\mathsf{parallel},\mathsf{prime},\mathsf{linear}\}$ and << is the order $\bigcup_{t\in\gamma^{-1}(\mathsf{linear})}<_{t}$ which ensures that tournaments are consistent with the corresponding $_{t}$. Notice that this labeled and partially-ordered tree is now uniquely determined from $G$. Moreover, edges from $F$ that connect children of a node not labeled by $\mathsf{prime}$ can be recovered from the so-refined weakly-partitive tree. In particular, if no nodes of $T$ is labeled by $\mathsf{prime}$, $(T,F)$ and thus $G$ is fully characterized by $(T,\gamma,<)$. Graphs for which this property holds are known as directed cographs. Directed cographs equivalently can be defined by a finite set of forbidden subgraphs (see [CP06]) and in the undirected case, cographs are exactly the $P_{4}$-free graphs. Directed cographs can be described by the refined weakly-partitive tree, called cotree, explained above and formalized in the following statement.
Theorem 18.
@@
Let $G$ be a directed cograph and let $T$ be the laminar tree induced by the family of its strong modules. There exists a unique total labeling $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{series},\mathsf{parallel},\mathsf{linear}\}$ of labels, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a unique linear ordering $_{t}$ of its children, such that every inner node having exactly two children is labeled $\mathsf{series}$ or $\mathsf{parallel}$, and the following condition is satisfied:
@@
  • for every two leaves $x$ and $y$ of $T$, denoting by $t$ their least common ancestor and $s_{x},s_{y}$ the children of $t$ that are ancestors of $x,y$ respectively, $xy$ is an edge of $G$ if and only if either $t$ is labeled by $\mathsf{linear}$ and $s_{x}<_{t}s_{y}$, or $t$ is labeled by $\mathsf{series}$.
@@
We naturally model cotrees as $\{\mathsf{ancestor},\mathsf{series},\mathsf{ord}\}$-structures as follow. A cotree $(T,\gamma,<)$ is modeled by $\mathbb{C}$ where $U_{\mathbb{C}}=V(T)$, $\mathsf{series}_{\mathbb{C}}=\gamma^{-1}(\mathsf{series})$, and $\mathsf{ord}_{\mathbb{C}}=\{(x,y)\mid x<y\}$. The nodes that are labeled by $\mathsf{linear}$ could be recovered as those inner nodes whose children are related by $\mathsf{ord}$, while the nodes that are labeled by $\mathsf{parallel}$ could be recovered as those inner nodes which are labeled neither by $\mathsf{series}$ nor by $\mathsf{linear}$. Based on 18 and as a consequence of 17, we can design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which produces the cotree of a cograph $G$ from $G$.
Corollary 19.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each directed cograph $G$ modeled by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{ancestor},\mathsf{series},\mathsf{ord}\}$-structure $\mathbb{C}$ representing the cotree $(T,\gamma,<)$ of $G$.
As for undirected graphs the set system of modules is partitive, a simpler result in which the label $\mathsf{linear}$ is not needed holds for undirected cographs.
5. Transducing split and bi-join decompositions